Ammonia is an important industrial raw material that is produced industrially by the Haber process according to this equation: N2+3H2 -> 2NH3. How many moles of NH3 can be formed from 250moles of N2 and 600 moles of H2? (N=14, H=1).


Deadline for assignment: 07/04/2022 (12pm)

1
Expert's answer
2022-04-06T15:50:04-0400

N2+3H2 = 2NH3

According to the equation, n (NH3) = 2 x n (N2) = 2/3 x n (H2).

Given amount of reactants can result in the following amount of NH3:

n (NH3) = 2 x n (N2) = 2 x 250 = 500 mol

n (NH3) = 2/3 x n (H2) = 2/3 x 600 = 400 mol

Therefore, H2 is the limiting reactant resulting in 400mol of NH3.